ELECTRONIC STATE OF CoH: A PROBLEM RELATED TO THE SIZE-INCONSISTENCY OF THE MR-SDCI METHOD AND THE MIXING OF ^{3}Φ AND ^{3}Π

\underline{REI FUKUI}, UMPEI NAGASHIMA, and TSUNEO HIRANO, Grid Technology Research Center, Institute of Advanced Industrial Science and Technology, 6-9-3 Ueno, Taito-ku, Tokyo 110-0015, Japan; HIROSHI TATEWAKI, Computation Center and Institute of Natural Sciences, Nagoya City University, Nagoya 467-8501, Japan; TAKESHI NORO AND KIYOSHI TANAKA, Division of Chemistry, Graduate School of Science, Hokkaido University, Kita 10, Nishi 8, Kita-ku, Sapporo, 060-0810, Japan.

The ground state of CoH has been determined experimentally as the ^{3}Φ by the far-infrared magnetic resonance study,^{a} and theoretically by ab initio methods.^{b} However, MR-SDCI and MR-SDCI + Q ab initio molecular orbital calculations predicted that the lowest electronic state is ^{5}Φ. The size-consistent MR-ACPF method correctly predicted that ^{3}Φ should be the ground state. As is previously reported by Tanaka et al.^{c} on FeH, the controversial results in MR-SDCI may arise through truncation error inherent in the size-inconsistent MR-SDCI method. Another difficulty is that we cannot clearly separate the Φ and Π states when CoH is treated under C_{2v} symmetry, not under its real symmetry of C_{∞v}. The best results we obtained at the moment are: r_{e} = 1.528 Å, B_{e} = 7.291 cm^{−1}, and ω_{e} = 2005 cm^{−1} at the level of the MR-ACPF + relativistic correction using the Roos ANO (Co) and Dunning aug-cc-pVQZ (H) basis sets.

aS.P. Beaton, K.M. Evenson, T. Nelis, and J.M. Brown, J. Chem. Phys., 89, 4446-4448 (1988).

bD.P. Chong, S.R. Langhoff, C.W. Bauschlicher, Jr., S.P. Walch, and H. Partridge, J. Chem. Phys., 85, 2850-2860 (1986);

M. Freindorf, C. M. Marian, and B. A. Hess, J. Chem. Phys., 99, 1215-1223 (1993).

cK. Tanaka, M. Sekiya, and M. Yoshimine, J. Chem. Phys., 115, 4558-4564 (2001).